When you purchased your car, you took out a five-year annual-payment loan with an interest rate of 6% per year. The annual payment on the car is $5,000. You have just made a payment and have now decided to pay off the loan by repaying the outstanding balance. What is the payoff amount for the following scenarios?


a) You have owned the car for one year (so there are four years left on the loan).

b) You have owned the car for four years (so there is one year left on the loan).
Answer #1

Annual payment of loan = $5000

Periodic interest rate (annual) = 6%

no of periods (n) = 5

First, We will be calculating loan amount :

(1+r) Payment = P*r* (1+r) - 1

Where, P= Loan amount

r = Periodic Interest rate = 0.06

n= no of periods = 5

Payment = $ 5000

(1 +0.06) 5000 = P* 0.06 * (1 + 0.06)5 - 1

5000 = P*0.06 * 1.338226 1.338226 - 1

Loan amount = $ 21061.84

a). Calculating the outstanding balance amount of loan if you have owned the car for one year:

Outstanding Balance = P* [(1 + r) - (1+r) (1 + r) - 1

Where, P= Loan amount = $ 21061.84

r = Periodic Interest rate = 0.06

n= no of periods = 5

m = no of lapsed = 1

[(1 +0.06) – (1 + 0.061 Outstanding Balance = 21061.84 * 1 (1 +0.06)5 – 1

1.338226 – 1.06 Outstanding Balance = 21061.84 * 1 1.338226-1

Outstanding balance of loan after 1 year = $ 17325.55

b) Calculating the outstanding balance amount of loan if you have owned the car for four year:

Outstanding Balance = P* [(1 + r) - (1+r) (1 + r) - 1

Where, P= Loan amount = $ 21061.84

r = Periodic Interest rate = 0.06

n= no of periods = 5

m = no of lapsed = 4

[(1 +0.06) – (1 +0.0641 Outstanding Balance = 21061.84 * 1 (1 +0.06)5 – 1

Outstanding Balance = 21061.84* 1.338226 - 1.262477 1.338226 - 1

Outstanding balance of loan after 4 year = $ 4717

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating